SOLVE CUBIC EQUATION 15X^3 + 11X^2 - 12X = 0
Answer by stanbon(75887)

15X^3 + 11X^2 - 12X = 0
X(15x^2+11x-12)=0
x(5x-3)(3x+4)=0
x=0 or x=3/5 or x=-4/3
Cheers,
Stan H.